The accompanying table shows results of challenged referee calls in a major tennis tournament. Use a 0.05 significance levelMore Info Was the Challenge to the Call Successful? Yes No 314 997 208 947 Men Women Print Done
Answer #1

I ran the tests again and every calculation and the values do match. I think the problem is cause the fields are demanding 4 decimal places and you're inputting 3 decimal place answers. So i'll add the fourth decimal in the images below so to make them more accurate. Check the final values and re enter.
The following cross tablulation have been provided. The row and column total have been calculated and they are shown below: Y
Squared Distances Yes No Men (314-277.511) 277.511 4.798 (997-1033.4892 1033.489 1.288 Women (208-244.4892 244.489 = 5.446 (9
Test Statistics The Chi-Squared statistic is computed as follows: n (Oij – Eij) =Σ Eij 4.798 + 5.446 + 1.288 + 1.462 = 12.994